Suppose x has a distribution with μ = 25 and σ = 23.
(a) If a random sample of size n = 44 is drawn, find μx, σ x and P(25 ≤ x ≤ 27). (Round σx to two decimal places and the probability to four decimal places.)
μx =
σ x =
P(25 ≤ x ≤ 27) =
(b) If a random sample of size n = 74 is drawn, find μx, σ x and P(25 ≤ x ≤ 27). (Round σ x to two decimal places and the probability to four decimal places.)
μx =
σ x =
P(25 ≤ x ≤ 27) =

Answer by stanbon(75887)

Suppose x has a distribution with μ = 25 and σ = 23.
(a) If a random sample of size n = 44 is drawn, find μx-bar, σ x-bar and
P(25 ≤ x-bar ≤ 27). (Round σx=bar to two decimal places and the probability to four decimal places.)
μx-bar = ux = 25
σ x-bar = 23/sqrt(44) = 3.4674
-----
z(25) = 0
z(27) = (25-27)/3.4674 = 0.5768
P(25 ≤ x-bar ≤ 27) = P(0< z < 0.4768) = 0.1832
------------------------------------------------
(b) If a random sample of size n = 74 is drawn, find μx, σ x and P(25 ≤ x ≤ 27). (Round σ x to two decimal places and the probability to four decimal places.)
I'll leave this to you.
μx =
σ x =
P(25 ≤ x ≤ 27) =
Cheers,
Stan H.